OpenStudy (anonymous):

Don't Want Answer!!! Just want someone to help me, explain how to figure this out. How many solutions does the system of equations have? y-6x=-3 and 4y-24x=-16 a.one b.two c.infinetley many d. none

OpenStudy (loser66):

time the first one by 4, what do you have?

OpenStudy (anonymous):

I got 4y-24x=-12

OpenStudy (loser66):

isolate 4y from both equations , what do you have?

OpenStudy (anonymous):

-24x=-16 and -24x=-12

OpenStudy (loser66):

no, no, "isolate" means 4y is in one side, and others are in other side Like this: for the firs one 4y = -12+24x do the same with the second one, please

OpenStudy (anonymous):

oh, sorry. :P 4y=-16+24x

OpenStudy (loser66):

that is the first one, the second one?

OpenStudy (anonymous):

4y=-16-24x

OpenStudy (loser66):

Oh I am sorry, your first one is correct. Now I rewrite them 4(y -6x)=4*(-3) ------> 4y -24x =-12 isolate 4y --> 4y = -12 +24x 4y -24x=-16 isolate 4y ---> 4y = -16 +24x got me so far??

OpenStudy (anonymous):

yeah

OpenStudy (loser66):

you see, both them have the left sides are 4y, that means the right side are equal -12+24x = -16 +24x got this part?

OpenStudy (anonymous):

yup

OpenStudy (loser66):

for this, -24x both sides, what do you have?

OpenStudy (loser66):

don't worry about the weird answer, just do

OpenStudy (anonymous):

-12=-16

OpenStudy (loser66):

yeah!!! -12 =-16 . Does it make sense to you?

OpenStudy (anonymous):

yeah! The answer is D. Right? Because -12 can never be equal to -16.

OpenStudy (loser66):

yup , yuop uyoup yyyyyyyyyyyyup, you got it by yourself, I didn't say anything about the answer, hehehee CONGRAT

OpenStudy (anonymous):

yay! \('.')/ Thanks again!

OpenStudy (loser66):

:)